OpenStudy (anonymous):

When AG = 16 ft, find the area of the region that is NOT shaded. Round to the nearest tenth. Choices: 666.9 ft2 200.9 ft2 731.2 ft2 136.9 ft2

OpenStudy (anonymous):

Here's a picture of the circle.

OpenStudy (anonymous):

Doe's anyone know how to set up the problem with an equation?

OpenStudy (triciaal):

AG = GB = 16 1/2 *16*16 = area of triangle AGB

OpenStudy (tkhunny):

AG is a radius of the circle. \(\triangle AGB\) is a Right Triangle. It is also isosceles. Lots of stuff to know about it.

OpenStudy (anonymous):

I get the part with the area.

OpenStudy (anonymous):

I don't know about the triangle part.

OpenStudy (anonymous):

Is it 136.9 ft2?

zepdrix (zepdrix):

Here is one silly idea you could try :) Subtract the `area of this square` from the `area of the circle`, and then take one fourth of that! \[\Large\rm A=\frac{1}{4}\left(A_{\circ}-A_{\square}\right)\]Just an idea ^^

OpenStudy (triciaal):

this is only one approach Area of complete circle = pi*r^2 and r = AG = 16 Angle AGB = 90° Sector AB = 1/4 of circle area of sector not shaded = 3/4* pi*r^2 + 1/2 *b*h for this problem r = b = h = 16 so area not shaded =

zepdrix (zepdrix):

Ooo did I steal your thunder? >.< My bad...

OpenStudy (triciaal):

@zepdrix that would give the area shaded

zepdrix (zepdrix):

Oh not shaded :) lol my bad.

OpenStudy (triciaal):

@FacePlant the area not shaded = area of circle - area shaded

zepdrix (zepdrix):

\[\Large\rm A_{shaded}=\frac{1}{4}\left(A_{\circ}-A_{\square}\right)\]\[\Large\rm A=A_{\circ}-A_{shaded}\] Ok I think I fixed it ;d Too many A's though lol

OpenStudy (anonymous):

I'm still kind of confused. ;/

OpenStudy (triciaal):

@FacePlant to find the answer to the nearest tenth calculate to 2 decimal places then round up to 1 decimal place.

OpenStudy (triciaal):

area not shaded = 3/4 the circle + triangle this will be the same as area of circle - area of shaded region

OpenStudy (triciaal):

what are you still confused about? ask specific let me try and make it clearer

OpenStudy (anonymous):

You think you could give me step by step on how to start it?

OpenStudy (anonymous):

Sorry if i'm taking kind of long. :/

OpenStudy (triciaal):

do you see that area not shaded = 3/4 the circle + triangle this will be the same as area of circle - area of shaded region

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I think i remember it now.

OpenStudy (anonymous):

I will set up the equation.

OpenStudy (triciaal):

now choose one way and then evaluate I did the part circle and add triangle while zepdrix choose the whole circle and minus the shaded should have the same answer whichever one you use.

OpenStudy (anonymous):

I got it, and thank you so much! :)

OpenStudy (anonymous):

The answer is 200.9 ft2, right?

OpenStudy (triciaal):

no problem

OpenStudy (triciaal):

let me do it

OpenStudy (triciaal):

i got 731

OpenStudy (tkhunny):

You may have been told to use a specific value for \(\pi\)?

OpenStudy (tkhunny):

\(\dfrac{3}{4}\pi 16^2 + \dfrac{1}{2}16^{2} = 256\cdot\dfrac{3\pi + 2}{4} = 256\cdot 2.85619449 = 731.18578949\)